OpenStudy (anonymous):

does the square root of -3 exist or is it imaginary?

OpenStudy (adi3):

wht do u think

OpenStudy (alekos):

its imaginary \[\sqrt{-3} = \sqrt{-1}\sqrt{3} = i \sqrt{3}\]

OpenStudy (unklerhaukus):

imaginary numbers exist